# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v2

@@ 4.2 Ablation Study
To study the impact of preprocessing, we compare the resulting detection AUROCs of AnomalyDINO with and without preprocessing in Figure 3. Across four different configurations of AnomalyDINO, we observe that by incorporating the proposed preprocessing, the resulting AUROCs increase by approximately $2\%$ without significantly affecting the inference time. The effect of proper preprocessing seems to increase with higher resolutions. Additionally, we compare the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’, and find that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 9 and 10. Depending on the product category, suitable preprocessing steps lead to substantial improvements. And while in principle augmenting the reference sample with rotations may negatively impact the detection performance, we observe this in only very few examples. For the full discussion, see Section C.1.
Aggregation choice
@@
We also evaluated the proposed framework using Vision Transformers pre-trained on ImageNet (see Figure 3). While these backbones slightly outperform PatchCore, they give weaker features compared to DINOv2 and are incompatible with the proposed masking procedure (zero-shot masking does not function effectively). This underlines that DINOv2 is indeed excellently suited for visual AD. In addition, ViT-based architectures like DINOv2 easily allow handling images with varying resolutions. In this context, we find that operating at a resolution of 448 gives the best trade-off between performance and inference time, but even higher detection AUROCs are possible at higher resolutions.
In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L). The full results are given in Section C.3, Figure 14. We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k\geq 8$). Interestingly, larger architectures do not necessarily translate to higher performance. The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes. In contrast, the larger architecture sizes perform better on VisA, which often involves multiple and more complex objects. This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
@@ 5 Conclusion
This paper proposes a vision-only approach for one- and few-shot anomaly detection. Our method is based on similarities between patch representations extracted by DINOv2 [29]. We carefully design means to populate the nominal memory bank with diverse and relevant features while minimizing false positives utilizing zero-shot segmentation and simple data augmentation. Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution. Our pipeline caters to these requirements through its simplicity and computational efficiency. The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches. Thus, our approach achieves the best of both worlds: improved performance and reduced inference time, especially compared to more complex vision-language methods. Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
@@
The specific pipeline exemplified in the paper focuses on simplicity and high throughput. However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives. For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods like [14] to obtain the anomaly map from patch features. It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors. We also plan to improve the batched zero-shot performance of AnomalyDINO.
Broader Impacts
@@ Appendix A Detailed experimental results
In this section, we present additional visualization and results for the proposed method.
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 4: Examples – MVTec-AD. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings. The color coding is normalized by the max. score over ‘good’ test samples.
Further anomaly maps, predicted by AnomalyDINO, are presented in Figures 4 and 5. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6, respectively.
@@
query GT 1-shot 8-shot query GT 1-shot 8-shot
@@
The results presented in Tables 5 and 6 show that a) anomalies in some categories are more difficult to than others, b) detection and segmentation performance correlate strongly (measured in image-level AUROC and PRO), and c), that the performance of AnomalyDINO increases across the board with more available reference samples. In particular, more complex objects like ‘PCB3’ and ‘PCB4’ in VisA or ‘Transistor’ in MVTec-AD seem to benefit the most from more available reference samples. In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s).
Dataset Resolution Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
MVTec-AD 448 99.3 98.8 99.7 97.9 61.8 93.9
672 99.5 99.0 99.8 98.2 64.3 95.0
VisA 448 97.2 93.7 97.6 98.7 50.5 95.0
672 97.6 94.5 98.0 98.8 53.8 96.1
The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this. We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
Finally, we also report the full-shot performance, see Table 7. As the diversity within $\mathcal{M}$ is sufficiently high given the large number of reference samples, we only apply masing here (no augmentations). The results demonstrate that the performance nicely improves for all considered metrics.
@@ B.1 Semantic Anomalies
Figure 6: Example of a semantic anomaly in category ‘Cable’ (MVTec-AD). Depicted are a nominal reference sample, an example of a sensory anomaly (Figure 6(b)) and a semantic anomaly (Figure 6(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
@@ B.2 The importance of informative reference samples
The second failure case occurs if the reference sample(s) does not resemble all concepts of normality, therefore $\mathcal{M}$ does not capture all variations of the nominal distribution $p_{\mathrm{norm}}$. As an illustration, consider the nominal samples of ‘Capsule’ in MVTec-AD, which may be rotated in such a way that the imprinted text is hidden (see Figure 7(b)). Therefore, parts of the text of a nominal test sample may be falsely recognized as anomalies. Due to the strong dependency on a suitable reference sample, we observe higher AD variances for some products in the one-shot setting, as shown in Tables 5 and 6. We like to remark, that this is only relevant to the few-shot setting, and with an increasing number of reference samples (and higher diversity of nominal patches in $\mathcal{M}$), the variance decreases notably.
@@ C.1 Preprocessing
As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ and preprocess test samples, influence the detection performance. The results per object for MVTec-AD and VisA are given in Figures 9 and 10, respectively.
@@
The same holds—although to a lesser extent—for the categories ‘Hazelnut’, ’Cable’ and ‘Wood’ (all MVTec-AD). However, we also observe that rotations of the reference sample can also decrease the detection performance in some categories, namely ‘PCB1/2/3’ and ‘Macaroni1’ (all VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 5, right column). And rotations of nominal samples may falsely reduce the distance between those patches depicting bent connectors and the nominal memory bank $\mathcal{M}$. Such a failure case based on a suboptimal preprocessing decision is depicted in Figure 11. The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 11). This highlights the importance of carefully designing the pre- and postprocessing pipeline for each object/product considered.
@@
Figure 12: Masking examples from MVTec-AD for all categories that passed the masking test (on a single train sample, see Table 8). Examples have not been cherry-picked.
@@
Figure 13: Masking examples from VisA (not cherry-picked).
@@
The results are given in Table 9. The average inference time per sample with AnomalyDINO-S (448) amounts to approximately 60ms in the few-shot regime ($\approx$ 16.7 samples/s), and only moderately increases with a larger memory bank (to 67ms ($+11\%$) for the full-shot scenario without augmentations). Compared to SOTA competitors in the one-shot regime, this is at least one order of magnitude faster (see Figure 3). Without any preprocessing steps, the 1-shot inference time amounts to approximately 55ms per sample ($\approx$ 18 samples/s). When applying both masking and rotations, the runtime increases from 55ms to 67ms, an moderate increase of roughly 23% (compared to the scenario without any preprocessing steps). The informed scenario can reduce the time to build the memory bank, but inference time is only affected for larger sample sizes. As expected, higher resolutions and larger architectures lead to increased runtimes.
@@ C.2 Scoring
In Section 3 we investigate different ways of aggregating the anomaly scores $\mathcal{D}$ on patch-level (in our case, distances to the nominal memory bank $\mathcal{M}$) to an image-level anomaly score via a statistic $q$. Note that the segmentation results are therefore not affected by the choice of $q$.
A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow [33] and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$. We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and the maximum of $\mathcal{D}$.
@@ C.3 Architecture Size
(a) Detection ARUOC for MVTec-AD.
(b) Detection ARUOC for VisA.
Figure 14: Effect of model size on detection AUROC for MVTec-AD and VisA. We report mean and standard deviation for the 1-shot setting over three seeds. The image resolution of AnomalyDINO is set to 672. Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by [20].
@@ Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
We need to infer anomaly scores for each patch representation $\mathbf{p}_{\mathrm{test}}$ of $\mathbf{x}^{(j)}$. We could again assess the distances between $\mathbf{p}_{\mathrm{test}}$ and $\mathcal{M}_{j}$ based on the distance to the nearest neighbor, as done in Equation 2. Note, however, that $\mathcal{M}_{j}$ may now contain nominal and anomalous patches. Thus, the nearest-neighbor approach is not suitable anymore: the nearest neighbor in $\mathcal{M}_{j}$ for (the representation of) an anomalous patch might also be abnormal, and the resulting distance therefore not informative.

### v3

@@ 4.2 Ablation Study
To study the impact of preprocessing, we compare the resulting detection AUROCs of AnomalyDINO with and without preprocessing in Figure 3. Across four different configurations of AnomalyDINO, we observe that by incorporating the proposed preprocessing, the resulting AUROCs increase by approximately $2\%$ without significantly affecting the inference time. The effect of proper preprocessing seems to increase with higher resolutions. Additionally, we compare the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’, and find that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 11 and 12. Depending on the product category, suitable preprocessing steps lead to substantial improvements. And while in principle augmenting the reference sample with rotations may negatively impact the detection performance, we observe this in only very few examples. For the full discussion, see Section C.1.
Aggregation statistic
@@
We also evaluated the proposed framework to Vision Transformers pre-trained on ImageNet, see Figure 3 and Sec. C.4. While these backbones slightly outperform PatchCore, they give weaker features compared to DINOv2 and are incompatible with the proposed masking procedure. This underlines that DINOv2 is indeed excellently suited for visual AD. In addition, ViT-based architectures like DINOv2 easily allow handling images with varying resolutions. In this context, we find that operating at a resolution of 448 gives the best trade-off between performance and inference time, but even higher detection AUROCs are possible at higher resolutions. We also evaluated our pipeline with DINOv2 at different distillation sizes (ViT-S, ViT-B, ViT-L). The full results are given in Section C.3, Figure 16. We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k\geq 8$). Interestingly, larger architectures do not necessarily translate to higher performance. The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes. In contrast, the larger architecture sizes perform better on VisA, which often involves multiple and more complex objects. This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
@@ 5 Conclusion
This paper proposes a vision-only approach for one- and few-shot anomaly detection. Our method is based on similarities between patch representations extracted by DINOv2 [30]. We carefully design means to populate the nominal memory bank with diverse and relevant features while minimizing false positives by utilizing zero-shot segmentation and simple data augmentation. Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution. Our pipeline caters to these requirements through its simplicity and computational efficiency. The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches. Thus, our approach achieves the best of both worlds: improved performance and reduced inference time, especially compared to more complex vision-language methods. Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
@@
The specific pipeline exemplified in the paper focuses on simplicity and high throughput. However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives. For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods like [15]. It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors. We also plan to improve the batched zero-shot performance of AnomalyDINO.
Acknowledgement
The authors acknowledge funding from TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport by the German Research Foundation (DFG).
Supplementary Material
@@ Appendix A Detailed experimental results
Query GT 1-shot 8-shot
Figure 4: Examples – MVTec-AD (1/2). Depicted are, from left to right, a test sample per category (Query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings. The color coding is normalized by the max. score over ‘good’ test samples.
Further anomaly maps, predicted by AnomalyDINO, are presented in Figures 4, 5, 6 and 7. The full results per category for MVTec-AD and VisA are given in Tables 6 and 7, respectively.
The results presented in Tables 6 and 7 show that a) anomalies in some categories are more difficult to than others, b), that the performance of AnomalyDINO increases across the board with more available reference samples. In particular, more complex objects like ‘PCB3’ and ‘PCB4’ in VisA or ‘Transistor’ in MVTec-AD seem to benefit the most from more available reference samples. In addition, we see that also the variance in the reported metrics decreases with the number of nominal samples.
In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s). [^6] Note that this holds for all one- and few-shot methods, not only for AnomalyDINO. The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this. We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
Finally, we also report the full-shot performance, see Table 5. As the diversity within $\mathcal{M}$ is sufficiently high given the large number of reference samples, we only apply masing here (no augmentations). The results demonstrate that the performance further improves for all considered metrics. Notably, AnomalyDINO-S (672) achieves new state-of-the-art segmentation performance measured in (AU)PRO in the full-shot setting (see here, accessed 11/26/2024).
Dataset Resolution Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
MVTec-AD 448 99.3 98.8 99.7 97.9 61.8 93.9
672 99.5 99.0 99.8 98.2 64.3 95.0
VisA 448 97.2 93.7 97.6 98.7 50.5 95.0
672 97.6 94.5 98.0 98.8 53.8 96.1
Query GT 1-shot 8-shot
Figure 5: Examples – MVTec-AD (2/2). See Fig. 4 for a description.
@@
Query GT 1-shot 8-shot
@@
Query GT 1-shot 8-shot
Figure 7: Examples – VisA (2/2). See Fig. 6 for the description.
@@ B.1 Semantic Anomalies
Figure 8: Example of a semantic anomaly in category ‘Cable’ (MVTec-AD). Depicted are a nominal reference sample, a sensory anomaly (Figure 8(b)), and a semantic anomaly (Figure 8(c)) with anomaly maps predicted by AnomalyDINO (1-shot).
@@ B.2 The importance of informative reference samples
The second failure case occurs if the reference sample(s) does not resemble all concepts of normality, therefore $\mathcal{M}$ does not capture all variations of the nominal distribution $p_{\mathrm{norm}}$. As an illustration, consider the nominal samples of ‘Capsule’ in MVTec-AD, which may be rotated in such a way that the imprinted text is hidden (see Figure 9(b)). Therefore, parts of the text of a nominal test sample may be falsely recognized as anomalies. Due to the strong dependency on a suitable reference sample, we observe higher AD variances for some products in the one-shot setting, as shown in Tables 6 and 7. We like to remark, that this is only relevant to the few-shot setting, and with an increasing number of reference samples (and higher diversity of nominal patches in $\mathcal{M}$), the variance decreases notably.
@@ C.1 Preprocessing
As discussed in Section 3 (and in the previous paragraph), suitable means to fill $\mathcal{M}$ and preprocess test samples, influence the detection performance. The results per object for MVTec-AD and VisA are given in Figures 11 and 12, respectively.
@@
The same holds—although to a lesser extent—for the categories ‘Hazelnut’, ’Cable’, and ‘Wood’ (all MVTec-AD). However, we also observe that rotations of the reference sample can also decrease the detection performance in some categories, namely ‘PCB1/2/3’ and ‘Macaroni1’ (all VisA), and to a small extent also ‘Transistor’ (MVTec-AD).
Figure 14: Masking examples from MVTec-AD for all categories that passed the masking test (on a single train sample, see Table 8).
We attribute this to the fact the specific anomalies for these printed circuit boards contain rotated or bent connectors (see the PCB examples in Figure 6, right column). And rotations of nominal samples may falsely reduce the distance between those patches depicting bent connectors and the nominal memory bank $\mathcal{M}$.
Such a failure case based on a suboptimal preprocessing decision is depicted in Figure 13. The anomaly refers to a rotated transistor (anomaly type ‘misplaced’), and when (falsely) rotating the reference sample, such anomalies will not be detected—in contrast to the ‘informed’ preprocessing (right side of Figure 13). This highlights the importance of carefully designing the pre- and postprocessing pipeline for each object/product considered.
Figure 15: Masking examples from VisA.
@@
The results are given in Table 9. The average inference time per sample with AnomalyDINO-S (448) amounts to approximately 60ms in the few-shot regime ($\approx$ 16.7 samples/s), and only moderately increases with a larger memory bank (to 67ms ($+11\%$) for the full-shot scenario without augmentations). Compared to SOTA competitors in the one-shot regime, this is at least one order of magnitude faster (see Figure 3). Without any preprocessing steps, the 1-shot inference time amounts to approximately 55ms per sample ($\approx$ 18 samples/s). When applying both masking and rotations, the runtime increases from 55ms to 67ms, a moderate increase of roughly 23% (compared to the scenario without any preprocessing steps). The informed scenario can reduce the time to build the memory bank, but inference time is only affected for larger sample sizes. As expected, higher resolutions and larger architectures lead to increased runtimes.
@@ C.2 Scoring
In Section 3 we investigate different ways of aggregating the anomaly scores $\mathcal{D}$ on patch-level (in our case, distances to the nominal memory bank $\mathcal{M}$) to an image-level anomaly score via a statistic $q$. Note that the segmentation results are therefore not affected by the choice of $q$. A standard choice is upsampling the patch distances of lower resolution to the full resolution of the test image using bilinear interpolation, then applying Gaussian smoothing operation (we follow [34] and set $\sigma=4.0$), to obtain an anomaly map $\mathcal{A}$, and set $q=\max{(\mathcal{A})}$. We also evaluate two potential alternatives of $q$, our default choice $q=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ (mean of the 1% highest entries in $\mathcal{D}$) and $q=\max(\mathcal{D})$.
@@ C.3 Architecture Size
(a) Detection AUROC for MVTec-AD.
(b) Detection AUROC for VisA.
Figure 16: Effect of model size on detection AUROC for MVTec-AD and VisA (mean and std over three seeds). The image resolution of AnomalyDINO is set to 672. Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation [21].
@@
(a) Detection AUROC for MVTec-AD.
(b) Detection AUROC for VisA.
Figure 17: Effect of backbone choice DINOv2 (trained in a self-supervised fashion) compared to ViT (supervised training on ImageNet) on detection AUROC for MVTec-AD and VisA (mean and std over three seeds). For better comparability, image resolution of AnomalyDINO is 224 for all models.
C.4 Backbone Choice
In our experiments, we find that DINOv2 provides excellently suited features for few-shot AD. This is already evident in Fig. 3, here we investigate the effect on the detection performance on MVTec-AD and VisA in more detail.
We compare the performance in few-shot AD of ViT-B and ViT-L pre-trained on ImageNet to that of AnomalyDINO-B and AnomalyDINO-L (utilizing DINOv2 in distillation sizes B and L). To ensure comparability, we deactivate any pre- and postprocessing for AnomalyDINO (no augmentations, no zero-shot masking) and set the image resolution to 224 for all models.
The results are depicted in Figure 17. We observe that DINOv2 significantly improves the detection performance compared to those of the ViT pre-trained on the image classification tasks, giving a performance gain of at least $+4\%$ AUROC on MVTec-AD and VisA. This demonstrates that the features extracted by DINOv2 are better suited compared to those of ImageNet-pretrained ViT-{B/L}.
In Sections 3 and 4 we demonstrate that further (substantial) improvements are possible with suitable pre- and postprocessing like augmentations and zero-shot masking (which is not possible based on the features from supervised features) and that the image resolution (or the effective patch size) can also greatly boost performance. See also Fig. 16 for the performance in dependence of the model size.
@@ Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
We need to infer anomaly scores for each patch representation $\mathbf{p}_{\mathrm{test}}$ of $\mathbf{x}^{(j)}$. We could again assess the distances between $\mathbf{p}_{\mathrm{test}}$ and $\mathcal{M}_{j}$ based on the distance to the nearest neighbor, as done in Equation 2. Note, however, that $\mathcal{M}_{j}$ may now contain nominal and anomalous patches. Thus, the nearest neighbor approach is not suitable anymore: the nearest neighbor in $\mathcal{M}_{j}$ for (the representation of) an anomalous patch might also be abnormal, and the resulting distance therefore not informative.